# Ordered random walks and the Airy line ensemble

### v1

@@ 4 Lower bound on the harmonic function
It follows from the definition of $h$ that $h(x)\geq\Delta(x)$. We will show now that on $x\in W_{t,\varepsilon}$ these functions are of the same order.
@@
Suppose $1/2-\varepsilon>2a$. There exists $t_{0}>0$ and $c>0$ such that for $t>t_{0},\gamma\in(0,1/2-2a-\varepsilon)$ and $x\in W_{t,\varepsilon}$,
@@
Since $\mathbf{E}e^{u\zeta}\leq 1+Cu$ for some constant $C$ and all sufficiently small $u$ we further obtain
@@
Let $p\geq 1$. Then, there exists a finite $C$ such that for $x\in W$ satisfying $x_{i+1}-x_{i}\geq 1$ for $11\leq i\leq d-$ 1,
@@
$\mathbf{E}_{x}[|\Delta(S(t)|^{p}])^{1/p}$ $\leq\Delta(x)\left(\mathbf{E}\left[\prod_{i<j}\left(1+|S_{j}(t)-S_{i}(t)|\right)^{p}\right]\right)^{1/p}.$
@@
$\mathbf{E}\left[\prod_{i<j}\left(1+|S_{j}(t)-S_{i}(t)|\right)^{p}\right]$ $\leq\mathbf{E}\left[\prod_{i<j}\left(1+|S_{j}(t)|+|S_{i}(t)|\right)^{p}\right]$
$\leq\mathbf{E}\left[\prod_{i<j}\left((1+|S_{j}(t)|)(1+|S_{i}(t)|)\right)^{p}\right]$ $\leq\mathbf{E}(1+|S_{1}(t)|)^{d^{2}p}\leq Ct^{d^{2}p},$
where we make use of the exponential inequality $\mathbf{P}(|S_{1}(t)|>u)\leq e^{-ru}$ for $u>t$ in the last step. ∎
@@
Let $d(t)\leq t^{a}$ for some $2a<\frac{1}{2}-\varepsilon$. There exist $\gamma>0,C>0$ and $t_{0}>0$ such that for $t>t_{0}$ and $x\in W_{t,\varepsilon}$,
@@
The first inequality follows from $h(x)\geq\Delta(x)$ for $x\in W$. The third inequality follows from Lemma 10. The rest of the proof is devoted to the second inequality.
Fix $\delta>0$ which we will define later. Let $X^{\prime}_{ij}$ be independent random variables distributed as $X_{ij}$ conditioned on $\{|X_{ij}|\leq t^{\delta}\}$, let $S^{\prime}_{j}(t)$ be the corresponding random walks and $\tau^{\prime}$ their exit time from the Weyl chamber. Let $n_{t}(j)=\sup\{n\geq 0:\sum_{i=1}^{n}e_{ij}\leq t\}$.
@@
We have
$\mathbf{E}_{x}[\Delta(S(t));\tau>t]$
$\geq\mathbf{E}_{x}[\Delta(S(t));\tau>t,\max_{1\leq i\leq 2t,1\leq j\leq d(t)}|X_{ij}|\leq t^{\delta},\max_{j}n_{t}(j)\leq 2t]$
$=\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t,\max_{j}n_{t}(j)\leq 2t]\mathbf{P}\left(\max_{1\leq i\leq 2t,1\leq j\leq d(t)}|X_{ij}|\leq t^{\delta}\right)$
$\geq\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t]\mathbf{P}\left(\max_{1\leq i\leq 2t,1\leq j\leq d(t)}|X_{ij}|\leq t^{\delta}\right)$
$\hskip 28.45274pt-\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t,\max_{j}n_{t}(j)>2t].$
We will first estimate the last summand using the Hölder inequality and Lemma 11 with $p=2$
$\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t,\max_{j}n_{t}(j)>2t]$ $\leq\mathbf{E}_{x}[\Delta(S^{\prime}(t))^{2}]^{1/2}\mathbf{P}(\max_{j}n_{t}(j)>2t)^{1/2}$
$\leq C\Delta(x)t^{d^{2}}e^{-t}.$
Thus it remains to bound the first summand. Note that $u(x)$ is harmonic for the free random walk as a difference of harmonic functions. Also note that
@@
and $T=\min(T_{1},\ldots,T_{i-1})$. Observe that, by the optional stopping theorem,
$\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t]\geq\mathbf{E}_{x}[\Delta(S^{\prime}(t));T>t]=\Delta(x)-\mathbf{E}_{x}[\Delta(S^{\prime}(T));T\leq t].$
@@
Note that $u(S(t\wedge T))$ is a martingale. Hence,
@@
$\mathbf{E}_{x}[\Delta^{\prime}(S(T));T\leq t]\leq t^{\delta}\Delta(x)\left(\left(1+\frac{1}{t^{1/2-\varepsilon}}\right)^{\frac{d(d-1)}{2}}-1\right)\leq\frac{\Delta(x)}{t^{\gamma}}t^{\delta}$
for some $\gamma>0$. As a result,
$\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t]\geq\Delta(x)(1-t^{\delta-\gamma}).$
@@
$\mathbf{P}\left(\max_{1\leq i\leq 2t,1\leq j\leq d(t)}|X_{ij}|\leq t^{\delta}\right)=(1-\mathbf{P}(|X_{ij}|>t^{\delta}))^{2td(t)}\geq 1-\frac{1}{t}$
provided $|X_{ij}|$ has sufficiently many moments. ∎
@@
Fix $\delta\in(0,1/2)$. There exist $C>0$, $t_{0}>1$ such that uniformly in $x\in W,d=d(t)$ and $t>t_{0}$,
@@
Fix $A$, which we will define later and set $b_{t}=[At^{1/2-\delta}]$. If $x\in W_{t,\delta}$, then there is nothing to prove since $\nu_{t}=0$. Hence, we will assume that $x\in W\cap W_{t,\delta}^{c}$. Next note that $W\cap W^{c}_{t,\delta}=\cup_{j=1}^{d-1}W_{j,t,\delta}$, where
$W_{j,t,\delta}=\left\{x\in W\colon\min_{1\leq k<j}(x_{k+1}-x_{k})>t^{1/2-\delta},x_{j+1}-x_{j}\leq t^{1/2-\delta}\right\}.$
@@
where $H(x)$ is the renewal function of descending ladder height process of the compound Poisson process $S_{j+1}(s)-S_{j}(s)$. Note that $H$ is the same for all $j$. Hence, by subadditivity of the renewal function we obtain the following uniform bound for $x\in W\cap W_{t,\delta}$,
@@
since $S_{t^{1-\delta}-b_{t}^{2}}\in W\cap W_{t,\delta}^{c}$ on the event $\{\tau>t^{1-\delta}-b_{t}^{2},\nu_{t}>t^{1-\delta}-b_{t}^{2}\}$. Repeating the procedure $\frac{t}{b_{t}^{2}}-1$ times we obtain,
@@
Let $x\in W_{{T},\varepsilon}$. Fix some $\delta\in(\varepsilon,1/2)$ and set $t=T^{\frac{1/2-\varepsilon}{1/2-\delta}}$. Then $x\in W_{t,\delta}$ and, in particular, Lemma 12 is applicable with $\varepsilon=\delta$.
@@
Note that by Lemma 10,
@@
$\mathbf{E}_{x}[\Delta(S({t_{k}}));\tau>t_{k},\nu_{t_{k+1}}>t_{k}]\leq\mathbf{E}_{x}[|\Delta(S({t_{k}})|^{p}])^{1/p}\mathbf{P}_{x}(\tau>t_{k},\nu_{t_{k+1}}>t_{k})^{1/q}$
for $11/p+1/q=$ 1. Then, by Lemma 11
$\mathbf{E}_{x}[|\Delta(S({t_{k}})|^{p}])^{1/p}\leq Ct_{k}^{d^{2}}.$
Applying Lemma 13 where $\varepsilon$ in Lemma 13 is chosen to be $\delta$ here.
@@
Now recall that $V(x)=\lim_{t\to\infty}\mathbf{E}_{x}[\Delta(S(t);\tau>t]$. Hence, repeating the above iterations
@@
Finally Lemma 12 shows that for $x\in W_{T,\varepsilon}$,
@@ 5.2 Preliminary lemmas
Let $\varepsilon^{\prime}$ be such that $\varepsilon^{\prime}<\frac{1}{2}-2a$. Choose $\delta^{\prime}>0$ sufficiently small.
For sufficiently small $c>0$ there exists $U,W\subset\mathbb{R}$ such that $P(X\in U)>c$, $P(X\in W)>c$ and $\sup\{x:x\in U\}<\inf\{x:x\in W\}$. For each of the embedded discrete-time random walks $S_{j}$ we either accept/reject each proposed step respectively. For $i\geq 1$, if $X_{ij}\in U$ we reject the proposed step with probability $c(j-1)/((d-1)\mathbf{P}(X_{ij}\in U))$ and if $X_{ij}\in W$ we reject with probability $c(d-j)/((d-1)\mathbf{P}(X_{ij}\in W))$. Otherwise the proposed step is accepted. Let $A_{t}$ denote the event that the $X_{ij}$ are accepted for all $1\leq i\leq dt^{1/2-\varepsilon^{\prime}+\delta^{\prime}}$ and $j=1,\ldots,d$. Note that $\mathbf{P}(A_{t})=p^{d^{2}t^{1/2-\varepsilon^{\prime}+\delta^{\prime}}}$ for $p:=1-c$. Let $X_{ij}^{\prime}$ denote $X_{ij}$ conditioned on the step being accepted and $S_{j}^{\prime}$ the corresponding random walks.
@@
Consider the event $D_{t}$ where the $d$-th walker jumps $dt^{1/2-\varepsilon^{\prime}+\delta^{\prime}}$ times while the other walkers stay in the same place. Then after the $k$-th walker has performed its jumps the $(k-1)$-th walker jumps $dt^{1/2-\varepsilon^{\prime}+\delta^{\prime}}$ times while the other walkers do not move, and so on. Note that
@@
Moreover, since the conditioned random walks $S^{\prime}$ satisfy (12)
@@
Applying Doob’s exponential inequality we can see now that both sums converge to 00 proving (5.2) for sufficiently large $t$.
@@
$\geq\mathbf{E}_{x}[V(S(\sigma_{t}));\tau>\sigma_{t},\sigma_{t}\leq t,A_{t},D_{t}]$
$\geq\mathbf{E}_{x}[V(S^{\prime}(\sigma_{t}));\sigma_{t}\leq t,\tau>\sigma_{t},S^{\prime}(\sigma_{t})\in W_{t,\varepsilon^{\prime}},|D_{t}]\mathbf{P}(A_{t}\cap D_{t}).$
@@
We will make use of the following proposition that couples continuous-time random walks with Brownian motion. In the discrete-time case this follows from Sakhanenko [36, Theorem 1].
@@
There exists $t_{0}>0$ such that for all $t\geq t_{0}$ there exists a Brownian motion $B_{j}$ such that
@@
For $j=1,\ldots,d$ set $U_{j}(n)=\sum_{i=1}^{n}X_{ij}$ for $n\geq 1$ and let $N_{j}(t)$ be a standard Poisson process. Then we have the following representation
@@
Next note that $(S_{j}(n))_{n\geq 0}$ is a random walk with increments distributed as $\sum_{i=1}^{N_{j}(1)}X_{ij}$. Hence, by Sakhanenko [36, Theorem 1] there exists a Brownian motion $B_{j}$ such that for $\delta>0$
@@
Let $\mathcal{C}_{t}:=\cap_{j=1}^{d}\mathcal{C}_{j,t}$, where
$\mathcal{C}_{j,t}:=\left\{\sup_{0\leq s\leq t}|B_{j}(s)-S^{\prime}_{j}(s)|\leq t^{\beta}\right\}.$
@@
    Suppose $2a<1/2-\varepsilon$. There exists $t_{0}>0$ and $C>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$, for all $x\in W_{t,\varepsilon}$ with $\lvert x_{j}\rvert\ll t^{\theta}$,
@@
    There exists $t_{0}>0$ and $C>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$, for all $x\in W^{d}$ with $\lvert x_{j}\rvert\ll t^{\theta}$ and $x_{j}-x_{j-1}\geq 1$ for all $j=2,\ldots,d,$
    $\mathbf{E}_{x}[\Delta(B(t));\tau^{bm}>t,\mathcal{C}_{t}^{c}]$ $\leq C\Delta(x)e^{\theta t^{a}\log(2t)-t^{\beta/2}}.$
@@
and note that
$\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{t}^{c}]\leq\sum_{j=1}^{d}\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t]$
$+\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{T}_{t}^{c}]+\mathbf{E}_{x}[V(S(t));\tau>t,\eta_{d}>t,\mathcal{T}_{t}].$ (15)
If $\eta_{d}\leq t$ and $S_{d}(t)-S_{1}(t)\leq t^{\theta}$ then $h(S(t))\leq(2t)^{\theta d}h_{[j]}(S(t))$. Therefore
$\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t]$ $\leq\mathbf{E}_{x}[h(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t]$
@@
where we used the fact that $h_{[j]}(S_{[j]}(t))1(\tau_{[j]}>t)$ is a supermartingale. Now applying Proposition 15
@@
$\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t]$ $\leq Ch_{[j]}(x)\exp\left(\theta t^{a}\log(2t)-t^{\beta/2}\right)$

### v2

@@ 4 Lower bound on the harmonic function
It follows from the definition of $h$ that $h(x)\geq\Delta(x)$ for $x\in W^{d}$. We will show now that on $x\in W_{t,\varepsilon}$ these functions are of the same order.
@@
Suppose $a\in(0,1/4)$ and $\varepsilon>0$ satisfy that $1/2-\varepsilon>2a$. There exists $t_{0}>0$ and $c>0$ such that for $t>t_{0},\gamma\in(0,1/2-2a-\varepsilon)$ and $x\in W_{t,\varepsilon}$,
@@
Since $\mathbf{E}[e^{u\zeta}]\leq 1+Cu$ for some constant $C$ and all sufficiently small $u$ we further obtain
@@
Let $p\geq 1$. There exists $t_{0}>0$ and a finite $C$ such that for all $t>t_{0}$ and $x\in W^{d}$ satisfying $x_{i+1}-x_{i}\geq 1$ for $1\leq i\leq d-1$,
@@
$(\mathbf{E}_{x}[|\Delta(S(t))|^{p}])^{1/p}$ $\leq\Delta(x)\left(\mathbf{E}_{0}\left[\prod_{i<j}\left(1+|S_{j}(t)-S_{i}(t)|\right)^{p}\right]\right)^{1/p}.$
@@
$\mathbf{E}_{0}\left[\prod_{i<j}\left(1+|S_{j}(t)-S_{i}(t)|\right)^{p}\right]$ $\leq\mathbf{E}_{0}\left[\prod_{i<j}\left(1+|S_{j}(t)|+|S_{i}(t)|\right)^{p}\right]$
$\leq\mathbf{E}_{0}\left[\prod_{i<j}\left((1+|S_{j}(t)|)(1+|S_{i}(t)|)\right)^{p}\right]$ $\leq\mathbf{E}_{0}(1+|S_{1}(t)|)^{d^{2}p}\leq Ct^{d^{2}p},$
where we make use of the following exponential inequality: there exists an $r>0$ such that $\mathbf{P}_{0}(|S_{1}(t)|>u)\leq e^{-ru}$ for $u>t$. The exact value of the exponent $r$ does not need to be chosen optimally and plays no role. ∎
@@
Let $a\in(0,1/4)$ and $\varepsilon>0$ with $d(t)\leq t^{a}$ for some $2a<\frac{1}{2}-\varepsilon$. There exist $\gamma>0,C>0$ and $t_{0}>0$ such that for $t>t_{0}$ and $x\in W_{t,\varepsilon}$,
@@
The first inequality follows from $h(x)\geq\Delta(x)$ for $x\in W^{d}$. The third inequality follows from Lemma 10. The rest of the proof is devoted to the second inequality.
Fix $\delta>0,$ which we will define later. Let $X^{\prime}_{ij}$ be independent random variables distributed as $X_{ij}$ conditioned on $\{|X_{ij}|\leq t^{\delta}\}$, let $S^{\prime}_{j}(t)$ be the corresponding random walks started from $x_{j}$ and $\tau^{\prime}$ their exit time from the Weyl chamber. Let $n_{t}(j)=\sup\{n\geq 0:\sum_{i=1}^{n}e_{ij}\leq t\}$ and $n_{t}=\sum_{j=1}^{d}n_{t}(j)$ be the total number of jumps up to time $t$. Let $(\widehat{X}_{k})_{k\geq 1}$ be a relabelling of the $X_{ij}$ according to the time at which the corresponding jump occurs, so that $(\widehat{X}_{k})_{1\leq k\leq n_{t}}$ is the collection of jumps that happen before time $t$.
We have
$\mathbf{E}_{x}[\Delta(S(t));\tau>t]$
$\geq\mathbf{E}_{x}[\Delta(S(t));\tau>t,\max_{1\leq k\leq 2d(t)t}|\widehat{X}_{k}|\leq t^{\delta},n_{t}\leq 2d(t)t]$
$=\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t,n_{t}\leq 2d(t)t]\mathbf{P}\left(\max_{1\leq k\leq 2d(t)t}|\widehat{X}_{k}|\leq t^{\delta}\right)$
$\geq\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t]\mathbf{P}\left(\max_{1\leq k\leq 2d(t)t}|\widehat{X}_{k}|\leq t^{\delta}\right)$
$\hskip 28.45274pt-\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t,n_{t}>2d(t)t].$
We will first estimate the last summand using the Hölder inequality. Then we use Lemma 11 with $p=2$ along with the fact that $n_{t}\sim\text{Po}(d(t)t)$ and applying a Chernoff bound to obtain
$\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t,n_{t}>2d(t)t]$ $\leq\mathbf{E}_{x}[\Delta(S^{\prime}(t))^{2}]^{1/2}\mathbf{P}(n_{t}>2d(t)t)^{1/2}$
$\leq C\Delta(x)t^{d^{2}}e^{-d(t)t/2}.$
@@
Note that
@@
and $T=\min(T_{1},\ldots,T_{d-1})$. Applying the optional stopping theorem to the stopping time $T\wedge t$, we have
$\Delta(x)$ $=\mathbf{E}[\Delta(S^{\prime}(T\wedge t))]$
$=\mathbf{E}[\Delta(S^{\prime}(t));T>t]+\mathbf{E}[\Delta(S^{\prime}(T));T\leq t].$
Combining this with the inequality $\tau^{\prime}\geq T$, we obtain
$\mathbf{E}[\Delta(S^{\prime}(t));\tau^{\prime}>t]\geq\mathbf{E}[\Delta(S^{\prime}(t));T>t]=\Delta(x)-\mathbf{E}[\Delta(S^{\prime}(T));T\leq t].$
@@
Note that $\Delta(x^{+})$ is harmonic for the free random walk using the same argument applied to $h$ in Proposition 8 (taking $\zeta_{j}=1$ for $j=1,\ldots,d$). Therefore, $u(x)$ is harmonic for the free random walk as a difference of harmonic functions and $u(S(t\wedge T))$ is a martingale. Hence,
@@
$\mathbf{E}_{x}[\Delta^{\prime}(S(T));T\leq t]\leq t^{\delta}\Delta(x)\left(\left(1+\frac{1}{t^{1/2-\varepsilon}}\right)^{\frac{d(d-1)}{2}}-1\right)\leq\frac{\Delta(x)}{t^{\gamma_{1}}}t^{\delta}$
for some $\gamma_{1}>0$. As a result,
$\mathbf{E}_{x}[\Delta(S^{\prime}(t));\tau^{\prime}>t]\geq\Delta(x)(1-t^{\delta-\gamma_{1}}).$
@@
$\mathbf{P}\left(\max_{1\leq k\leq 2td(t)}|\widehat{X}_{k}|\leq t^{\delta}\right)=(1-\mathbf{P}(|\widehat{X}_{k}|>t^{\delta}))^{2td(t)}\geq 1-\frac{1}{t}$
provided $|X_{ij}|$ and their relabelling as $\widehat{X}_{k}$ have sufficiently many moments, which is a consequence of our assumptions in Section 2.2. Finally, taking $\delta=\gamma_{1}/2$ and $\gamma=\gamma_{1}/2$ gives the required statement. ∎
@@
Fix $\delta\in(0,1/2)$. There exist $C>0$, $t_{0}>1$ such that uniformly in $x\in W^{d},d=d(t)$ and $t>t_{0}$,
@@
Fix $A$, which we will define later and set $b_{t}=At^{1/2-\delta}$. If $x\in W_{t,\delta}$, then there is nothing to prove since $\nu_{t}=0$. Hence, we will assume that $x\in W^{d}\cap W_{t,\delta}^{c}$. Next note that $W^{d}\cap W_{t,\delta}^{c}=\cup_{j=1}^{d-1}W_{j,t,\delta}$, where
$W_{j,t,\delta}=\left\{x\in W^{d}\colon\min_{1\leq k<j}(x_{k+1}-x_{k})>t^{1/2-\delta},x_{j+1}-x_{j}\leq t^{1/2-\delta}\right\}.$
@@
where $H(x)$ is the renewal function of descending ladder height process of the compound Poisson process $S_{j+1}(s)-S_{j}(s)$. Note that $H$ is the same for all $j$. Hence, by subadditivity of the renewal function we obtain the following uniform bound for $x\in W^{d}\cap W_{t,\delta}$,
@@
since $S_{t^{1-\delta}-b_{t}^{2}}\in W^{d}\cap W_{t,\delta}^{c}$ on the event $\{\tau>t^{1-\delta}-b_{t}^{2},\nu_{t}>t^{1-\delta}-b_{t}^{2}\}$. Repeating the procedure $\frac{t}{b_{t}^{2}}-1$ times we obtain,
@@
Let $x\in W_{{T},\varepsilon}$. Fix some $\delta\in(\varepsilon,1/2-2a)$ and set $t=T^{\frac{1/2-\varepsilon}{1/2-\delta}}$. Then $x\in W_{t,\delta}$ and, in particular, Lemma 12 is applicable with $\varepsilon=\delta$.
@@
Here, $c$ denotes a positive constant whose value can change from line to line. Note that by Lemma 10,
@@
$\mathbf{E}_{x}[\Delta(S({t_{k}}));\tau>t_{k},\nu_{t_{k+1}}>t_{k}]\leq(\mathbf{E}_{x}[|\Delta(S({t_{k}}))|^{p}])^{1/p}\mathbf{P}_{x}(\tau>t_{k},\nu_{t_{k+1}}>t_{k})^{1/q}$
for $1/p+1/q=1$. Then, by Lemma 11, taking $p=q=2$,
$(\mathbf{E}_{x}[|\Delta(S({t_{k}}))|^{p}])^{1/p}\leq C\Delta(x)t_{k}^{d^{2}}.$
Applying Lemma 13 where $\varepsilon$ in Lemma 13 is chosen to be $\delta$ here,
@@
The condition that $2a<\delta/(1-\delta)$ ensures that the first time in this iteration, $k=0$, decays to zero. Now recall that $V(x)=\lim_{t\to\infty}\mathbf{E}_{x}[\Delta(S(t);\tau>t]$. Hence, repeating the above iterations
@@
Finally, Lemma 12 shows that for $x\in W_{T,\varepsilon}$,
@@ 5.2 Preliminary lemmas
In the proof, we will introduce a time parameter $t>0$ satisfying $d(t)\leq t^{a}$ for $a<1/4.$ Let $\varepsilon^{\prime}$ be such that $\varepsilon^{\prime}<\frac{1}{2}-2a$. Choose $\delta^{\prime}>0$ sufficiently small.
For sufficiently small $c>0$ there exists $U,W\subset\mathbb{R}$ such that $P(X\in U)>c$, $P(X\in W)>c$ and $\sup\{x:x\in U\}<\inf\{x:x\in W\}$. For each of the embedded discrete-time random walks $S_{j}$ we either accept/reject each proposed step respectively. For $i\geq 1$, if $X_{ij}\in U$ we reject the proposed step with probability $c(j-1)/((d-1)\mathbf{P}(X_{ij}\in U))$ and if $X_{ij}\in W$ we reject with probability $c(d-j)/((d-1)\mathbf{P}(X_{ij}\in W))$. Otherwise the proposed step is accepted. Let $A_{t}$ denote the event that the $X_{ij}$ are accepted for all $1\leq i\leq dt^{1/2-\varepsilon^{\prime}+\delta^{\prime}}$ and $j=1,\ldots,d$. Note that $\mathbf{P}(A_{t})=p^{d^{2}t^{1/2-\varepsilon^{\prime}+\delta^{\prime}}}$ for p:=1cp:=1-c. Let $X_{ij}^{\prime}$ denote $X_{ij}$ conditioned on the step being accepted and $S_{j}^{\prime}$ the corresponding random walks.
@@
To justify this, let $\mathcal{R}$ denote the event that a proposed step is rejected. Then,
$\mathbf{E}[X_{i,j+1};\mathcal{R}^{c}]-\mathbf{E}[X_{i,j};\mathcal{R}^{c}]=\mathbf{E}[X_{i,j};\mathcal{R}]-\mathbf{E}[X_{i,j+1};\mathcal{R}]$
$=\frac{c(j-1)\mathbf{E}[X_{ij};X_{ij}\in U]}{(d-1)\mathbf{P}(X_{ij}\in U)}+\frac{c(d-j)\mathbf{E}[X_{ij};X_{ij}\in W]}{(d-1)\mathbf{P}(X_{ij}\in W)}$
$\qquad-\frac{c(j+1-1)\mathbf{E}[X_{ij};X_{ij}\in U]}{(d-1)\mathbf{P}(X_{ij}\in U)}-\frac{c(d-j-1)\mathbf{E}[X_{ij};X_{ij}\in W]}{(d-1)\mathbf{P}(X_{ij}\in W)}$
$\geq\frac{c}{d-1}\left(\mathbf{E}[X_{ij}|X_{ij}\in W]-\mathbf{E}[X_{ij}|X_{ij}\in U]\right).$
Using $\sup\{x:x\in U\}<\inf\{x:x\in W\}$, we can obtain (13).
Consider the event $D_{t}$ where the $d$-th walker jumps $dt^{1/2-\varepsilon^{\prime}+\delta^{\prime}}$ times while the other walkers stay in the same place. Then after the $k$-th walker has performed its jumps the (k1)(k-1)-th walker jumps $dt^{1/2-\varepsilon^{\prime}+\delta^{\prime}}$ times while the other walkers do not move, and so on. Note that
@@
Moreover, since the conditioned random walks $S^{\prime}$ satisfy (13),
@@
Applying Doob’s exponential inequality we can see now that both sums converge to 0 proving (5.2) for sufficiently large $t$.
@@
$\geq\mathbf{E}_{x}[V(S(\sigma_{t}));\sigma_{t}\leq t,\tau>\sigma_{t},A_{t},D_{t}]$
$\geq\mathbf{E}_{x}[V(S^{\prime}(\sigma_{t}));\sigma_{t}\leq t,\tau>\sigma_{t},S^{\prime}(\sigma_{t})\in W_{t,\varepsilon^{\prime}}|D_{t}]\mathbf{P}(A_{t}\cap D_{t}).$
@@
We will make use of the following proposition that couples continuous-time random walks with Brownian motion. In the discrete-time case this follows from Sakhanenko [39, Theorem 1].
@@
Let $\beta>0.$ There exists $t_{0}>0$ such that for all $t\geq t_{0}$ there exists a Brownian motion $B_{j}$ such that
@@
For $j=1,\ldots,d$ set $U_{j}(n)=\sum_{i=1}^{n}X_{ij}$ for $n\geq 1$ and let $N_{j}(t)$ be a standard Poisson process. Then we have the following representation,
@@
Next note that $(S_{j}(n))_{n\geq 0}$ is a random walk with increments distributed as $\sum_{i=1}^{N_{j}(1)}X_{ij}$. Hence, by Sakhanenko [39, Theorem 1] there exists a Brownian motion $B_{j}$ such that for $\delta>0$
@@
Let $\mathcal{C}_{t}:=\cap_{j=1}^{d}\mathcal{C}_{j,t}$, where for $\beta>0$,
$\mathcal{C}_{j,t}:=\left\{\sup_{0\leq s\leq t}|B_{j}(s)-S_{j}(s)|\leq t^{\beta}\right\}.$
Recall $\tau^{\text{bm}}=\inf\{t\geq 0:B(t)\notin W^{d}\}$.
@@
    Suppose $\varepsilon\in(0,1/2)$ and $2a<1/2-\varepsilon$. There exists $t_{0}>0$ and $C>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$, for all $x\in W_{t,\varepsilon}$ with $\lvert x_{j}\rvert\leq t^{\theta}/\log(t)$,
@@
    There exists $t_{0}>0$ and $C>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$, for all $x\in W^{d}$ with $\lvert x_{j}\rvert\leq t^{\theta}/\log(t)$ and $x_{j}-x_{j-1}\geq 1$ for all $j=2,\ldots,d,$
    $\mathbf{E}_{x}[\Delta(B(t));\tau^{\mathrm{bm}}>t,\mathcal{C}_{t}^{c}]$ $\leq C\Delta(x)e^{\theta t^{a}\log(2t)-t^{\beta/2}}.$
@@
and recall the definition of $h$ in terms of $\eta_{1},\ldots,\eta_{d}$ from Section 3. Note that
$\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{t}^{c}]\leq\sum_{j=1}^{d}\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t^{\theta}]$
$+\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{T}_{t}^{c}]+\mathbf{E}_{x}[V(S(t));\tau>t,\eta_{d}>t^{\theta},\mathcal{T}_{t}].$ (16)
If $\eta_{d}\leq t^{\theta}$ and $S_{d}(t)-S_{1}(t)\leq t^{\theta}$ then $h(S(t))\leq(2t)^{\theta d}h_{[j]}(S(t))$. Therefore,
$\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t^{\theta}]$ $\leq\mathbf{E}_{x}[h(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t^{\theta}]$
@@
where we used the fact that $h_{[j]}(S_{[j]}(t))1(\tau_{[j]}>t)$ is a supermartingale. Now applying Proposition 15,
@@
$\mathbf{E}_{x}[V(S(t));\tau>t,\mathcal{C}_{j,t}^{c},\mathcal{T}_{t},\eta_{d}\leq t^{\theta}]$ $\leq Ch_{[j]}(x)\exp\left(\theta t^{a}\log(2t)-t^{\beta/2}\right)$